OpenStudy (anonymous):

find the absolute maximum and minimum for the function f(x)=x+yz on the ball B = {(x, y, z) | x^2 + y^2 + z^2 =< 1}.

OpenStudy (zarkon):

find all your critical points ....throw out any not in your ball...then use Lagrange multipliers to check the boundary condition

OpenStudy (anonymous):

I found that x=1, and y=z which imply that critical point is (1,0,0) how to find the minimum

OpenStudy (zarkon):

why is that the only critical point? Also you need to check the entire boundary

OpenStudy (anonymous):

How to find other critical point?

OpenStudy (amistre64):

f(x)=x+yz ?? or f(x,y,z)=x+yz ??

OpenStudy (anonymous):

yes, you are right f(x,y,z)=x+yz subject to x^2+y^2+z^2<=1 find maximium and minimum I use Lagrange multipliers and found that x=1 and y^2=z^2 by plug in into constraint, I found critical point (1,0,0) how to find other critical points?